# Language models should be subject to repeatable, open, domain-contextualized hallucination benchmarking

### v1

@@ Abstract
1 Introduction
@@
Language models should be subject to repeatable, open, domain-contextualized hallucination benchmarking. The current environment of hallucination benchmarking is hampering the development of models that could perform better and with fewer errors and hallucinations. Specifically, because benchmarks are (a) not open; (b) not repeatable; (c) are either domain ignorant or ablate expertize behind private walls, it is not possible for researchers and commercial developers to openly compare model performance in the pursuit of better functioning, safer, and less-carbon intensive frameworks. By developing open and repeatable benchmarks, the language modeling community can establish a shared set of tasks that support model training and experimentation prior to submitting models to public evaluation. When these benchmarks incorporate expert knowledge, they help ensure that resulting models are both valid and practically useful.
To identify and mitigate hallucinations in contemporary LLMs, a crucial first step is the development of a process to create a golden dataset that supports domain-specific hallucination benchmarks. One key lesson we learned—the hard way—is that domain expertise must be present from the inception of the dataset. Without domain expert involvement in the early stages of data creation, hallucination metrics lack scientific validity and real-world utility.
@@
This paper proceeds as follows: we begin by outlining the problem space of hallucination benchmarking, summarizing the state of the field, providing a working definition of model hallucination, and providing three examples the consequences of model hallucination in the real-world. We then argue why a repeatable, open, domain-contextualized approach will produce better outcomes. We then demonstrate how expert responses differ from model-generated text, drawing on one case study from medicine and a second from politics. We conclude by evaluating alternative perspectives.
2 Language Model Hallucination
Although there are several descriptive features about language models — the amount of time or carbon to train a model, the speed with which a model generates tokens at inference, inter alia — we focus specifically on controlling incorrect or misleading text, i.e., hallucination benchmarking. [^1] See Appendix B for a brief survey of model performance and accuracy. Recent advances in open and repeatable benchmarks (e.g., srivastava2023, ; bbeh2025, ) have helped to constrain general-purpose, model-generated text. The result is that LLM-generated text has become increasingly, though only seemingly, coherent. But the shine, has come off the stochastic parrot — simply generating seemingly coherent language is no longer sufficient, and language models are being pushed to produce outputs that serve concrete business, policy, medical and scientific purposes. Stakeholders increasingly want language models to take expert-like actions in high-stakes domains, but current models are not ready to operate at that level of rigor or reliability. Indeed, as models grow larger, the hallucination problem appears to be worsening (Zhou2024larger_and_more, ).
@@ 2 Language Model Hallucination
Early efforts in LLM hallucination benchmarking primarily centered on summarization tasks that compared model-generated summaries against reference texts (e.g., CNN and DailyMail (hermann2015, ; narayan2018, ) and XSum (narayan2018, )) to assess factual consistency. As LLMs evolved and were used to generate seemingly novel text, new benchmarks that did not rely upon scoring against reference texts became necessary. When benchmarking metrics scored only perceived coherence or helpfulness, teams’ trained models to score better on these measures, but with the side effect of wild hallucinations — entirely ungrounded references, links, and factual claims (see, for example, subsection 2.2) (huang2025survey, ). Recent advancements have moved from evaluating the perceived coherence of language, to assessing the factual accuracy of model-generated text in zero-shot question answering (see, e.g., Hughes_Vectara_Hallucination_Leaderboard_2023, ; li2023haluevallargescalehallucinationevaluation, ).
@@
We highlight two recent advancements in hallucination detection: the Vectara hallucination leaderboard and HaluEval. The Vectara hallucination leaderboard provides an open and continuously updated platform for evaluating hallucination rates across LLMs using document summary metric, the Hughes Hallucination Evaluation Model (HHEM) (Hughes_Vectara_Hallucination_Leaderboard_2023, ; hhem-2.1-open, ). HHEM assesses factuality using a zero-shot question-answering setup over grounded documents; humans define ground-truth, binary judgments of correctness. In a separate project, HaluEval collects 45,000 human-annotated samples across multiple generation tasks (e.g., question answering, summarization, open-domain generation), distinguishing between extractive, abstractive, and factual hallucinations (li2023haluevallargescalehallucinationevaluation, ). Crucially, HaluEval 2.0, provides a domain-relevant benchmarks across biomedicine, finance, science, education, and as well as an open category (li2024dawndarkempiricalstudy, ).
@@
Despite these advancements, we observe that the field has focused on only a small subset of possible dimensions from the full range available to evaluate and constrain models. These selected dimensions often include tasks like causal reasoning, geometric rotation, and calendar scheduling under conditions of incomplete information (see, e.g. bbeh2025, ). Certainly these are difficult tasks for machines — without the ability to intervene, data simply cannot possess the full compliment of counterfactual measurements to define a tractable loss function — but they are trivial for any human. It is unsurprising that model performance tends to degrade on zero-shot tasks that fall outside the scope of these evaluations. The use of ever larger models has resulted in improved performance on in-benchmark tasks but also increased hallucination on out-of-benchmark tasks (Zhou2024larger_and_more, ).
2.1 Defining LLM Hallucination
@@
While there is a considerable body of work surrounding language model hallucination, the field has not settled on a single operational definition of what, specifically constitutes a hallucination (see e.g., ziwei2023, ; huang2025survey, , for a survey of the field). The term, hallucination, arises in 2017 in the context of a data-to-text generation task (wiseman2017challenges, ), the concept features in abstractive summarization in 2020 (maynez2020faithfulness, ), and features prominently in Stochastic Parrots (bender2021dangers, ). To create a point of reference within this paper, we define language model hallucinations in the following way:
Definition 1
@@
A response that does not include qualifying language fails to convey the probabilistic nature of the output, instead presenting ideas as definitive facts. Responses that appropriately include qualifying language — for example, in medicine — might use phrases such as “Those symptoms could be a sign of [diagnosis] but they could also be signs of [alternative diagnosis],” or “It is important to consult with a healthcare provider…” to communicate uncertainty and encourage additional verification.
@@
The greatest trick language models ever pulled was convincing the world they were coherent. Current generation language models produce syntactically correct, seemingly coherent strings of text (bender2021dangers, ). In particular, the challenge this presents is that “…Our perception of natural language text, regardless of how it was generated, is mediated by our own linguistic competence and our predisposition to interpret communicative acts as conveying coherent meaning and intent” (bender2021dangers, , p. 616). [^2] The authors have first-hand language of this type of incoherent language: when checking results of a statistics test they administered, they asked an LLM to generate solutions for all the test questions. In one case, the answers relied on theorems that were not presented in the course; in another case, the language models produced text that engaged in a loop of reasoning without advancing the solution concept toward a conclusive statement. In other words, because the text sounds like language that a human produces, we interpret the text to have meaning and intent.
@@
LLM text that misstates a verifiable fact is perhaps the most obvious form of hallucination. Huang et. al. identity two types of fact-based hallucinations: (1) factual contradiction where text presents statements that can be verified with real-world information; and (2) factual fabrication where models make up facts to support statements, but these facts are inconsistent with verifiable real-world knowledge. In subsection 2.2 we develop extended examples of this type of factual fabrication. Presently, factual contradiction is undertaken by selecting only “high-quality” data sources (see Lucy et. al. for important context surrounding data filtering (lucy2024aboutme, )), Retrieval Augmented Generation (RAG), few-shot training and knowledge graph based approaches, and, recently agent-based approaches. The profusion of approaches belies the reality that this type of factual incorrectness is the easiest type of problem to identify, and so has been the locus of considerable development.
@@
Unverified responses that are those that either fail to cite a source, or cite an external source that is insufficient to corroborate the claim — whether to confirm or contradict it (ziwei2023, ) (huang2025survey, ). For example, if a user asks about eligibility requirements to vote in California, an LLM hallucination may either fail to cite a source altogether or cite one — such as a nonprofit organization or the California Secretary of State — where the source cited is insufficient to verify the validity of the LLM response.
2.2 Consequences of Language Model Hallucination
@@
Language model hallucination is not simply a problem to be solved so that models score better on benchmarks. These models power widely used consumer products, and when users read and act on hallucinated output, there is a real risk of harmful consequences. Because communication is so core to the human experience, syntactically correct language serves as a credibility hack (chomsky1976, ; christiansen2008, ), effectively jailbreaking our perceptive system to imbue what seems like coherence from a stochastic parrot (bender2021dangers, ). LLM models can hide their lack of coherence within the written word — LLMs produce text that impersonates the writing of experts, and are imbued with “expertise” by readers. When combined with human decision making which is well-known to be prone to using lower-cost heuristic evaluation (simon1955, ; simon1956, ; kahneman2011, ) — even among high-stakes decision makers with very strong financial and political incentives for decision accuracy (hafner2013, ) — this jailbreak creates a profound risk for misinterpretation, and evaluation of model-generated language (bender2021dangers, ). In this section, we introduce three examples, two of which we present to domain experts.
Voters Seeking Political Process Information
@@
The manner for how elections are conducted is constitutionally devolved to the states (Article 1, Section 4, US Constitution). Although federal law stipulates an common election day, states are largely responsible for determining the means and methods for how elections are run. Consequently, if a voter were to move from one state to another, it is possible that the election process might look quite different. The differences may include: which state agency handles registration; registration requirements and dates; availability of absentee or early voting; voter identification requirements; among others. In this context, it would be quite reasonable for a voter to query an LLM with a question like, “What documentation do I need to vote, and when does early voting start in [my location]?” Queries of this type generated documented instances of hallucination that would lead people to the wrong early voting date. These hallucinations persisted even when the model was instructed to describe the “click-path” that was used to navigate from an official government source to the specific resource used to produce an election date; the model generated text that traversed to a non-existent series of sites (hughes2025, ). For some period of time ChatGPT 3.5-x returned a caution to these political information queries, encouraging users to confirm information with their local elections agencies; however, these cautions are no longer present for users who are using ChatGPT 4o in April 2025.333OpenAI is also aware of the limitations of ChatGPT for expert users in the political space. In 2024, the company prohibited politicians and lobbyists from using ChatGPT for official campaign activities(NPR2024, ). There is also evidence that LLMs generated inaccurate and misleading responses for voters during the 2024 U.S. presidential primary (nyt2025, ).
@@
Medical diagnosis is a challenging task. Patients noisily introspect their own state, commonly exhibit social desirability bias, and frequently present with comorbid diseases that are caused by a complex relationship between genetics, lifestyle, and proximate causes. Front line clinicians — nurse practitioners and general practitioners — develop expertise that incorporates medical science and real-world clinical training and use this expertise to ask winnowing questions and to order informative tests to arrive at an manageable set of possible diagnoses. With this set, doctors propose medical interventions and observe how patients respond.
Compared to clinicians, language models are missing huge swaths of the diagnostic toolkit. Models do not have access to long clinical histories or family or social context. They do not have the ability to manipulate, either physically to check the severity of an ankle sprain, or through prescription to see whether a fever responds to acetaminophen. And, models cannot order labs or imaging to distinguish between hidden states that are unknowable to a patient. [^4] In spite of this, some may may use LLMs as complete substitutes for visits to the doctor’s office. The costs of asking language model are substantially lower than attending a clinic: language models do not activate the same stigma responses as human-to-human conversation; querying a language model is substantially easier and quicker than scheduling and traveling to an in-person clinical visit; and, for under-resourced populations with limited clinical access, language models may not be an approximation or substitute for clinical care, it my be the only option. It is little surprise, then, that model-based diagnostic accuracy is substantially lower than diagnostic accuracy of expert doctors (takita2025systematic, ). To the extent that existing work includes medical expertise in the creation and benchmarking of language models, they are largely “bolting-on” doctors’ expert review at the end of the diagnostic pipeline (see, e.g., wu2024llmsciterelevantmedical, ; kim2025, ).
AI in the Criminal Justice System
Technology has a long history of being used in the field of criminal justice. Before the current era of generative artificial intelligence, AI/ML tools were mostly relegated to tasks such as surveillance, event prediction and crime prevention. When applied to such use cases, these tools were repeatedly found to reinforce or enable new systemic biases, compromise the privacy rights of citizens, and in many cases simply proved to be ineffective at their stated purpose (buolamwini2018gender, ; raji2020saving, ; taboh2021clearview, ). A shift toward more advanced models has taken place in recent years. For example, image restoration models can be used on biometric recognition (norman2024investigation, ), and LLMs (like ChatGPT) are increasigly integrated into lawyers’ workflows (tan2023chatgpt, ; izzidien2024llm, ). The efficacy of both applications is limited. For example, in 2025, lawyers who used an LLM tool to generate legal briefs were disciplined in court after it was discovered that documentation submitted contained hallucinated case law (lawhallucination, ). The potential for scaled harm is clear: in cases where the hallucinated citation is not detected, the error is carried forward until it is discovered, setting the foundation for potentially catastrophic and cascading legal crisis.
3 Solution Space
@@
Language model developers should have repeatable benchmarks that allow for internal experimentation and evaluation prior to submitting models for public benchmarking. Without repeatable internal benchmarks, teams must repeatedly submit against high-visibility, public evaluation benchmarks (e.g., Hughes_Vectara_Hallucination_Leaderboard_2023, ; li2023haluevallargescalehallucinationevaluation, ; hong2024hallucinationsleaderboardopen, ; bbeh2025, ) creating perverse consequences for the advancement of the science of language models.
@@
When consumers choose model adoption based on highly-visible leader-boards, scientific and profit incentives are muddied. Recall, for example, that in 2015 a research team claimed a 10.2% relative improvement over the state of the art. As would come to light through several revisions to the paper and considerable public hand-wringing by the contest organizers (ilsvrc2015announcement, ), the research team had submitted 200 distinct models to the scoring dataset, more than all other entrants combined (simonite2015, ). If development teams do not have repeatable benchmarks that can be used while building and training models, and instead have to evaluate their work on public scoring benchmarks, this problem will assuredly continue to arise (wiggers2025crowdsourced, ). Accelerating this problem are additional selection pressures: developers want to work on competitive teams, and technical decision makers want to purchase an implement the best performing models.
Open
@@
Openness and repeatability are principal tenants of scientific investigation. As a field, computer science has a long history of building computing tools in the pursuit of this openness (e.g., knuth1984texbook, ). Not only does openness allow for the development and transmission of ideas between researchers, research groups and universities, but it also allows for open evaluation and repeatable execution of tests conducted against data (kluyver2016jupyter, ). The open science movement that makes compute environments, scientific code, and data widely available has become de rigueur for publication across many scientific communities, speeding discovery and error correction, and reducing malfeasance (christensen2019, ). Benchmarking openness, as we intend it, combines aspects of scientific openness and open source software (OSS).
@@
Consistent with scientific openness, benchmark openness should be open about the benchmark questions, the benchmark code, and the benchmark scores. This type of scientific openness supports the repeatable benchmarking position presented in the last subsection. When consumers are exposed to technology products that are mere days off of the scientific development floor, these products must necessarily contain aspects of scientific evaluation. Openness about how a model performs against different benchmarks is a minimum “safety evaluation” that is due to consumers who are going to use a product. Just as medication labels contain statements of side effects, model publishers should describe the expected behaviors and error rates so that users can make minimally informed decisions (mitchell2019model, ). Consistent with the OSS movement, open benchmarks allow for distributed contributions from domain-experts ensuring benchmarks have broad topical coverage. What is more, open contributions distribute the maintenance costs for this evaluation tooling.
@@
The state of the art benchmarking for LLMs evaluates models on tasks that are challenging, yet are very narrow. General reasoning benchmarks evaluate reasoning capabilities across a range of broad tasks such adjudicating boardgames rule questions (kazemi2024, ), complex geometry problems (kazemi2023, ) and causal reasoning (bbeh2025, ). Developers have chosen to drive models toward performance on these narrowly defined tasks with the consequence that the models hallucinate at higher rates in other zero- and few-shot prompts (Zhou2024larger_and_more, ; nyt2025, ).
Simply stated, it is not possible for teams to develop effective benchmark evaluations in areas where they are non-expert. These teams lack knowledge of the key dimensions to consider for a topic outside their expertise as well as the skill to design questions that would evaluate evaluate models on those dimensions. These teams are able to write benchmarks that are difficult in their domain — solvable by the average humans or skilled language model researchers — but they fall short in representing deep knowledge in other fields. The result is that the best current community-generated, contextualized evaluation frameworks focus on reasoning tasks that do not generalize well to novel, domain-specific tasks. When models are developed without human expertise as a part of the training, they hallucinate more frequently when asked for expert-level advice compared to prompts that do not require expertise (Zhou2024larger_and_more, ).
4 Case Study in Hallucination Dataset Development

### v2

# The Case for Repeatable, Open, and Expert-Grounded Hallucination Benchmarks in Large Language Models

@@ Abstract
Keywords: LLM Evaluation, Benchmarking Methodologies, GenAI Explainability, Hallucination
[…]
The Case for Repeatable, Open, and Expert-Grounded Hallucination Benchmarks in Large Language Models
Justin D. Norman, Michael U. Rivera, D. Alex Hughes
University of California, Berkeley, School of Information
{justin.norman, michaelrivera, dhughes}@berkeley.edu
Abstract content
1.  Introduction
@@
The current landscape of hallucination benchmarking is hampering progress toward the development of models that perform efficiently and produce fewer errors and hallucinations. Because existing benchmarks are (a) not repeatable, (b) not open, and (c) either domain-agnostic or isolate expertize behind private walls, researchers and commercial developers cannot openly collaborate and compare model performance in the pursuit of faster, more capable, and less-carbon intensive frameworks. Developing open, repeatable benchmarks would enable the language modeling community to establish a shared foundation for model training and experimentation prior to submitting models to public evaluation. When these benchmarks incorporate expert knowledge, they help ensure that resulting models are both valid and practically useful.
@@
This paper proceeds as follows. We begin with an outline of the problem space of hallucination benchmarking. Next, we summarize the state of the field, provide a working definition of model hallucination, and illustrate the real-world consequences of hallucination through two examples. We then argue that a repeatable, open, domain-contextualized approach will produce fewer hallucinations. Next, we demonstrate how expert responses differ from model-generated text, drawing on a case study from medicine and another from politics. We conclude by highlighting some remaining challenges in the field.
A crucial first step in identifying and mitigating hallucinations in contemporary Large language models (LLMs) is the development of a process to create a golden dataset that supports domain-specific hallucination benchmarks. One key lesson we learned—the hard way—is that domain expertise must be present from the very inception of the dataset. Without the involvement of domain experts in the early stages of data creation, hallucination metrics risk lacking both scientific validity and real-world utility.
2.  Language Model Hallucination
Advances in open and repeatable benchmarks (e.g., srivastava2023; bbeh2025) have helped to constrain general-purpose, model-generated text resulting in text that has become increasingly, though only seemingly, coherent. But, what was initially regarded as an impressive novelty is not being incorporated into academic, civil-society, and corporate processes. Language models are being pushed to produce outputs that serve concrete business, policy, medical and scientific purposes and stakeholders are increasingly relying on language models to take expert-like actions in high-stakes domains. This creates an acute problem: it is very unlikely that current models are ready to operate at the level of rigor or reliability that is expected from a domain expert. Although it is generally known that recent model vintages are producing hallucinations Zhou2024larger_and_more; kalai2025languagemodelshallucinate, the implications of the widespread implementation of these models is less widely discussed. Indeed, although there is remarkable enthusiasm to use language models, there is virtually zero understanding of how these models perform at expert and near-expert tasks.
@@ 2.  Language Model Hallucination
Early efforts in LLM hallucination benchmarking primarily centered on summarization tasks that compared model-generated summaries against reference texts to assess factual consistency (e.g., CNN and DailyMail (hermann2015; narayan2018) and XSum (narayan2018)). As LLMs evolved and were used to generate seemingly novel text, new benchmarks that did not rely upon scoring against reference texts became necessary. When benchmarking metrics scored only perceived coherence or helpfulness, teams’ trained models to score better on these measures, but with the side effect of wild hallucinations—entirely ungrounded references, links, and factual claims. Recent advancements have moved from evaluating the perceived coherence of language, to assessing the factual accuracy of model-generated text in zero-shot question answering (see, e.g., Hughes_Vectara_Hallucination_Leaderboard_2023; li2023haluevallargescalehallucinationevaluation).
@@
We highlight two recent advancements in hallucination detection: the Vectara hallucination leaderboard and HaluEval. The Vectara hallucination leaderboard provides an open and continuously updated platform for evaluating hallucination rates across LLMs using document summary metric, the Hughes Hallucination Evaluation Model (HHEM) (Hughes_Vectara_Hallucination_Leaderboard_2023; hhem-2.1-open). HHEM assesses factuality using a zero-shot question-answering setup over grounded documents; humans define ground-truth, binary judgments of correctness. In a separate project, HaluEval collects 45,000 human-annotated samples across multiple generation tasks (e.g., question answering, summarization, open-domain generation), distinguishing between extractive, abstractive, and factual hallucinations (li2023haluevallargescalehallucinationevaluation). Crucially, HaluEval 2.0, provides a domain-relevant benchmarks across biomedicine, finance, science, education, and as well as an open category (li2024dawndarkempiricalstudy).
@@
Despite these advancements, we remark that the field has focused on only a small subset of possible dimensions from the full range available to evaluate and constrain models. These selected dimensions include tasks like causal reasoning, geometric rotation, and calendar scheduling under conditions of incomplete information (see, e.g. bbeh2025). Certainly these are difficult tasks for machines that are trivial for humans; but current implementations do not focus on calendar scheduling tasks. Said simply, existing hallucination benchmarks are not informed by the type of expertise that is necessary to constrain models working in hard contexts, and so hallucination continues to be a challenge for models working on expert-level tasks.
2.1.  Defining LLM Hallucination
@@
While there is a considerable body of work surrounding language model hallucination, the field has not settled on a single operational definition of what, specifically constitutes a hallucination (see e.g., ziwei2023; huang2025survey, for a survey of the field). The term hallucination arises in 2017 in the context of a data-to-text generation task (wiseman2017challenges), features in abstractive summarization in 2020 (maynez2020faithfulness), and features prominently in Stochastic Parrots (bender2021dangers). To create a point of reference within this paper, we define language model hallucinations in the following way:
Definition.
@@
Qualifying language is language that conveys either the probabilistic nature of the output or the lack of uncertainty about the true state of the world. A response that does not include such language makes a seemingly authoritative statement or fails to acknowledge the existence of alternatives.
@@
Current generation language models produce syntactically correct, seemingly coherent strings of text (bender2021dangers). Indeed, it was the initial emergence of seeming coherence, paired with an engaging user-facing design that led to the eruption of popular attention to language models in 2022. Bender and Gebru coin the evocative term “stochastic parrots” to highlight the risks of this seemingly coherent language, writing, “Our perception of natural language text, regardless of how it was generated, is mediated by our own linguistic competence and our predisposition to interpret communicative acts as conveying coherent meaning and intent(bender2021dangers, p. 616). In other words, because model generated text reads like language that a human expert might produce, humans have a blind-spot that leads us to interpret the text as possessing meaning, intent, and as we argue, expertise. First generation, non-expert-derived benchmarks have largely solved issues of coherence and model generated text is nearly uniformly coherent on general-language tasks.
@@
LLM text that misstates a verifiable fact is currently the most crucial form of hallucination to address for near-expert performance. Working on document summarization, Huang et. al. identity two types of fact-based hallucinations: (1) intrinsic errors where text presents statements that can be verified against known corpora; and (2) extrinsic errors where models make up facts to support statements, but these facts are inconsistent with verifiable real-world knowledge. Importantly, when working on expert-like tasks, both intrinsic and extrinsic errors require expertise to evaluate whether model generated text is correct or hallucinated. Extrinsic errors can be controlled in two ways: either (a) creating tighter safeguards and guardrails that cause models to more readily default to “I don’t know” behavior; or (b) providing more expertise. Presently, factual contradiction is undertaken by selecting only “high-quality” data sources (see Lucy et. al. for important context surrounding data filtering (lucy2024aboutme)), Retrieval Augmented Generation (RAG), few-shot training and knowledge graph based approaches, and, recently agent-based approaches. The profusion of approaches belies the reality that this type of factual incorrectness is the most pressing current problem, and so has been the locus of considerable development.
@@
Unverified responses that are those that either fail to cite a source, or cite an external source that is insufficient to corroborate the claim — whether to confirm or contradict it (ziwei2023) (huang2025survey). For example, if a user asks about eligibility requirements to vote in California, an LLM hallucination may either fail to cite a source altogether or cite one — such as a nonprofit organization or the California Secretary of State — where the source cited is insufficient to verify the validity of the LLM response.
$2.$  Consequences of Language Model Hallucination
@@
Language model hallucination is not simply a problem to be solved so that models score better on benchmarks. These models power widely used consumer products, and when users read and act on hallucinated output, there is a real risk of harmful consequences. Because communication is so core to the human experience, syntactically correct language serves as a credibility hack (chomsky1976; christiansen2008), effectively jailbreaking our perceptive system to imbue what seems like coherence from a stochastic parrot (bender2021dangers). LLM models can hide their lack of coherence within the written word — LLMs produce text that impersonates the writing of experts, and are imbued with “expertise” by readers. When combined with human decision making which is well-known to be prone to using lower-cost heuristic evaluation (simon1955; simon1956; kahneman2011) — even among high-stakes decision makers with very strong financial and political incentives for decision accuracy (hafner2013) — this jailbreak creates a profound risk for misinterpretation, and evaluation of model-generated language (bender2021dangers). In this section, we introduce two examples which we present to domain experts.
Voters Seeking Information on Political Process
@@
The manner for how elections are conducted is constitutionally devolved to the states (Article 1, Section 4, US Constitution). Although federal law stipulates an common election day, states are largely responsible for determining the means and methods for how elections are run. Consequently, if a voter were to move from one state to another, it is possible that the election process might look quite different. The differences may include: which state agency handles registration; registration requirements and dates; availability of absentee or early voting; voter identification requirements; among others. In this context, it would be quite reasonable for a voter to query an LLM with a question like, “What documentation do I need to vote, and when does early voting start in [my location]?” Queries of this type generated documented instances of hallucination that would lead people to the wrong early voting date. These hallucinations persisted even when the model was instructed to describe the “click-path” that was used to navigate from an official government source to the specific resource used to produce an election date; the model generated text that traversed to a non-existent series of sites (hughes2025). For some period of time ChatGPT 3.5-x returned a caution to these political information queries, encouraging users to confirm information with their local elections agencies; however, these cautions are no longer present for users who are using ChatGPT 4o in April 2025.111OpenAI is also aware of the limitations of ChatGPT for expert users in the political space. In 2024, the company prohibited politicians and lobbyists from using ChatGPT for official campaign activities(NPR2024). There is also evidence that LLMs generated inaccurate and misleading responses for voters during the 2024 U.S. presidential primary (nyt2025).
@@
Medical diagnosis is a challenging task. Patients noisily introspect their own state, commonly exhibit social desirability bias, and frequently present with comorbid diseases that are caused by a complex relationship between genetics, lifestyle, and proximate causes. Front line clinicians develop expertise that incorporates medical science and real-world clinical training and use this expertise to ask winnowing questions and to order informative tests to arrive at an manageable set of possible diagnoses. With this set, doctors propose medical interventions and observe how patients respond.
Compared to clinicians, language models are missing huge swaths of the diagnostic toolkit. Models do not have access to long clinical histories or family or social context; nor do they have the ability to manipulate, either physically to check the severity of an ankle sprain, or through prescription to see whether a fever responds to acetaminophen; nor can models order labs or imaging to distinguish between hidden states that are unknowable to a patient. [^2] In spite of this, it is likely that people view LLMs as substitutes for visits to the doctor’s office. It is little surprise, then, that model-based diagnostic accuracy is substantially lower than diagnostic accuracy of expert doctors (takita2025systematic). To the extent that existing work includes medical expertise in the creation and benchmarking of language models, they are largely “bolting-on” doctors’ expert review at the end of the diagnostic pipeline (see, e.g., wu2024llmsciterelevantmedical; kim2025).
3.  Solution Space
@@
Table 1: Expert vs. Language Model Response Comparison. The yellow text denotes correct expert information, and the green text denotes hallucinated LLM content.
Field Content
Prompt I’ve got frequent nosebleeds, bruising easily, and fatigue. Do I have aplastic anemia?
Expert Answer No, you might have idiopathic thrombocytopenia, von Willebrand disease, malignancy
LLM Answer Aplastic anemia is a rare condition in which the body stops producing enough new blood cells. Symptoms can include frequent nosebleeds, easy bruising, fatigue, and other signs of low blood cell counts such as paleness and shortness of breath. However, these symptoms can also be caused by other conditions. It is important to see a healthcare provider for a proper evaluation and diagnosis.
Hallucination Yes - Factual Contradiction and Factual Generalization
Hallucination Metrics $FCD^{+}=3.57,\ FGR^{+}=1.0,\ SCD^{-}=1.22,\ THS^{-}=2.57$
Impact Expert explicitly rules out aplastic anemia and provides specific alternative diagnoses. LLM confirms aplastic anemia as possibility, missing critical differential diagnoses. METEOR+=0.10 shows poor alignment with expert guidance.
@@
Language model developers should provide repeatable benchmarks that allow for internal experimentation and evaluation prior to submitting models for public benchmarking. Without repeatable internal benchmarks, teams must repeatedly submit against high-visibility, public evaluation benchmarks (e.g., Hughes_Vectara_Hallucination_Leaderboard_2023; li2023haluevallargescalehallucinationevaluation; hong2024hallucinationsleaderboardopen; bbeh2025) creating perverse consequences for the advancement of the science of language models.
@@
When consumers choose model adoption based on highly-visible leader-boards, scientific and profit incentives are muddied. Recall, for example, that in 2015 a research team claimed a 10.2% relative improvement over the state of the art (ilsvrc2015announcement). As would come to light through several revisions to the paper and considerable public hand-wringing by the contest organizers , the research team had submitted 200 distinct models to the scoring dataset, more than all other entrants combined (simonite2015). If development teams do not have repeatable benchmarks that can be used while building and training models, and instead have to evaluate their work on public scoring benchmarks, this problem will assuredly continue to arise (wiggers2025crowdsourced). Accelerating this problem are additional selection pressures: developers want to work on competitive teams, and technical decision makers want to purchase an implement the best performing models.
Openness
@@
Trust and repeatability are principal tenants of scientific investigation. As a field, computer science has a long history of building computing tools in the pursuit of this openness (e.g., knuth1984texbook). Not only does openness allow for the development and transmission of ideas between researchers, but it also allows for open evaluation and repeatable execution of tests conducted against data (kluyver2016jupyter). The open science movement that makes compute environments, scientific code, and data widely available has become de rigueur for publication across many scientific communities, speeding discovery and error correction, and reducing malfeasance (christensen2019). Benchmarking openness, as we intend it, combines aspects of scientific openness and open source software (OSS).
These principles are evident when write tests for students as instructors or teachers: We ask teaching assistants or graduate student instructors who are working in that class to write practice tests and test questions. Crafting these test questions often involve reviewing and refining those draft questions before they are actually made to be student facing. Under this model, the experts are contributing to both the practice sets (i.e. the public development and the actual test (i.e. the private evaluation sets) and we are relying on them to keep from sharing those answers with students. Current benchmarking frameworks make the draft sets public and benchmarking sets private, but do not have a mechanism for open contributions to either of these sources of ground truth. What we propose is not to break the hidden nature of the benchmark evaluation set, but rather that experts can contribute to the efforts.
@@
Consistent with scientific openness, benchmark openness should be open about the benchmark questions, the benchmark code, and the benchmark scores. This type of scientific openness supports the repeatable benchmarking position presented in the last subsection. When consumers are exposed to technology products that are mere days off of the scientific development floor, these products must necessarily contain aspects of scientific evaluation. Openness about how a model performs against different benchmarks is a minimum “safety evaluation” that is due to consumers who are going to use a product. Just as medication labels contain statements of side effects, model publishers should describe the expected behaviors and error rates so that users can make minimally informed decisions (mitchell2019model). Consistent with the OSS movement, open benchmarks allow for distributed contributions from domain-experts ensuring benchmarks have broad topical coverage. What is more, open contributions distribute the maintenance costs for this evaluation tooling.
@@
The state of the art benchmarking for LLMs evaluates models on tasks that are challenging, yet are very narrow. General reasoning benchmarks evaluate reasoning capabilities across a range of broad tasks such adjudicating boardgames rule questions (kazemi2024), complex geometry problems (kazemi2023) and causal reasoning (bbeh2025). Developers have chosen to drive models toward performance on these narrowly defined tasks with the consequence that the models hallucinate at higher rates in other zero- and few-shot prompts (Zhou2024larger_and_more; nyt2025).
Simply stated, it is not possible for teams to develop effective benchmark evaluations in areas where they are non-expert. These teams lack knowledge of the key dimensions to consider for a topic outside their expertise as well as the skill to design questions that would evaluate evaluate models on those dimensions. These teams are able to write benchmarks that are difficult in their domain — solvable by the average humans or skilled language model researchers — but they fall short in representing deep knowledge in other fields. The result is that the best current community-generated, contextualized evaluation frameworks focus on reasoning tasks that do not generalize well to novel, domain-specific tasks. When models are developed without human expertise as a part of the training, they hallucinate more frequently when asked for expert-level advice compared to prompts that do not require expertise (Zhou2024larger_and_more).
4.  Case Study in Hallucination Dataset Development